# The Power of Iterative Filtering for Supervised Learning with (Heavy) Contamination

### v1

@@ C.1 Low-Degree Approximators
Our structural Theorem 4.2 can be combined with results from polynomial approximation theory in order to obtain end-to-end results on learning with nasty noise, including the results of Table 2.
@@
For the class of $k$-halfspace intersections, [74] showed that the $\epsilon$-approximate degree with respect to the standard Gaussian distribution $\mathcal{N}_{d}$ (see Definition 4.1) is $\ell(\epsilon)\leq O(\log(k)/\epsilon^{2})$ and [63] obtains the same result with respect to the uniform distribution $\operatorname{Unif}_{d}$ over the hypercube.
@@
For the class of degree-$k$ PTFs, the $\epsilon$-approximate degree with respect to the standard Gaussian was bounded by $\ell(\epsilon)\leq O(k^{2}/\epsilon^{2})$ in [60]. In [62], it was shown that $\ell(\epsilon)\leq\frac{k^{O(k^{2})}(\log 1/\epsilon)^{O(k\log k)}}{\epsilon^{2}}$ with respect to the uniform over {±$\{\pm 1\}^{d}$.
@@ C.2 Sandwiching Approximators
For the more challenging problem of learning with heavy contamination, we require the existence of sandwiching approximators (see Theorem 4.4), and show that arbitrary (non-sandwiching) approximators do not suffice (see Theorem 5.2).
@@ Appendix D Complexity of Learning with Adversarial Label Noise
The complexity of agnostic learning with respect to Gaussian marginals was shown to be characterized in terms of the $\mathcal{L}_{1}$ approximation degree for algorithms that fall in the statistical query (SQ) framework by [37], implying the lower bounds of Table 4 for intersections of halfspaces and polynomial threshold functions. For the uniform distribution on the hypercube, a similar characterization was given by [42]. Moreover, [4] showed an information-theoretic lower bound for learning monotone functions even in the realizable setting, implying the lower bound of Table 4 for monotone functions. For $\mathsf{AC}^{0}$ circuits, [65] provided a cryptographic lower bound.

### v2

@@ C.1 Low-Degree Approximators
Our structural Theorem˜4.2 can be combined with results from polynomial approximation theory in order to obtain end-to-end results on learning with nasty noise, including the results of Table˜2.
@@
For the class of $k$-halfspace intersections, [75] showed that the $\epsilon$-approximate degree with respect to the standard Gaussian distribution $\mathcal{N}_{d}$ (see Definition˜4.1) is $\ell(\epsilon)\leq O(\log(k)/\epsilon^{2})$ and [66] obtains the same result with respect to the uniform distribution $\operatorname{Unif}_{d}$ over the hypercube.
@@
For the class of degree-$k$ PTFs, the $\epsilon$-approximate degree with respect to the standard Gaussian was bounded by $\ell(\epsilon)\leq O(k^{2}/\epsilon^{2})$ in [64]. In [65], it was shown that $\ell(\epsilon)\leq\frac{k^{O(k^{2})}(\log 1/\epsilon)^{O(k\log k)}}{\epsilon^{2}}$ with respect to the uniform over $\{\pm 1\}^{d}$.
@@ C.2 Sandwiching Approximators
For the more challenging problem of learning with heavy contamination, we require the existence of sandwiching approximators (see Theorem˜4.4), and show that arbitrary (non-sandwiching) approximators do not suffice (see Theorem˜5.2).
@@ Appendix D Complexity of Learning with Adversarial Label Noise
The complexity of agnostic learning with respect to Gaussian marginals was shown to be characterized in terms of the $\mathcal{L}_{1}$ approximation degree for algorithms that fall in the statistical query (SQ) framework by [35], implying the lower bounds of Table˜4 for intersections of halfspaces and polynomial threshold functions. For the uniform distribution on the hypercube, a similar characterization was given by [22]. Moreover, [8] showed an information-theoretic lower bound for learning monotone functions even in the realizable setting, implying the lower bound of Table˜4 for monotone functions. For $\mathsf{AC}^{0}$ circuits, [72] provided a cryptographic lower bound.